OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5.

OpenStudy (ash2326):

x intercept of 6 , and y intercept of 5 we know standard equation for a line with x and y intercept a and b is \[ x/a+ y/b=1\] here a =6 and y =5 \[ x/6+ y/5=1\] in slope intercept form \[ 5x+6y=30\] \[ y=-5/6x+5\] here slope =-5/6 and y intercept=5